OpenStudy (anonymous):

Help please, trig hw Solve for θ in the interval 0 ≤ θ ≤ 360. 3tanθ-2 = tanθ

OpenStudy (anonymous):

You can graph the two function and solve for the angle

OpenStudy (anonymous):

or you can combine the tangent functions \[3tan(\theta) - tan(\theta) = 2\] solve for theta \[tan(\theta) = 1 -> arctan(tan(\theta)) = arctan(1) -> \theta = arctan(1)\]